OpenStudy (anonymous):

The quadratic 3x2 – 9x – 10 has what type of roots?

OpenStudy (campbell_st):

use the descriminant a = 3, b = -9 and c = -10 discriminant is \[\Delta = b^2 - 4ac\] substituting \[\Delta = (-9)^2 - 4\times3\times-10\] \[\Delta = 201\] the quadratic has 2 unequal irrational roots...